Two very narrow slits are spaced 1.80 mm apart and are placed 35.0 cm from a screen.

What is the distance between the first and second dark lines of the interference pattern when the slits are illuminated with coherent light with wavelength 550 nm?



1
Expert's answer
2022-02-28T10:38:37-0500

"x=(2\cdot2+1-2\cdot 1-1)\frac{\lambda l}{2d}=\frac{\lambda l}d=107~\mu m."